### v1

@@ Abstract
We introduce a versatile flexible-captioning vision-language model (VLM) capable of generating region-specific descriptions of varying lengths. The model, FlexCap, is trained to produce length-conditioned captions for input bounding boxes, and this allows control over the information density of its output, with descriptions ranging from concise object labels to detailed captions. To achieve this we create large-scale training datasets of image region descriptions of varying length, starting from captioned images.
This flexible-captioning capability has several valuable applications. First, FlexCap demonstrates superior performance in dense captioning tasks on the Visual Genome dataset. Second, a visual question answering (VQA) system can be built by employing FlexCap to generate localized descriptions as inputs to a large language model. The resulting system achieves state-of-the-art zero-shot performance on a number of VQA datasets. We also demonstrate a localize-then-describe approach with FlexCap can be better at open-ended object detection than a describe-then-localize approach with other VLMs. We highlight a novel characteristic of FlexCap, which is its ability to extract diverse visual information through prefix conditioning. Finally, we qualitatively demonstrate FlexCap’s broad applicability in tasks such as image labeling, object attribute recognition, and visual dialog. Project webpage: https://flex-cap.github.io.
@@ 1 Introduction
The groundbreaking success of large language models (LLMs) has led to the widespread adoption of this technology in a variety of applications. The introduction of vision-language models (VLMs) has further enhanced these capabilities by enabling reasoning over visual content in the form of question answering and visual dialog. Yet, the ideal representation for utilizing LLMs to enable visual applications is still an open question.
A popular and effective strategy is to directly provide visual features as input tokens to LLMs. Two successful systems, Flamingo [2] and BLIP [29, 28], have focused on utilizing latent representations produced by frozen visual backbones and integrating these with frozen LLMs. While Flamingo uses cross-attention to adapt visual tokens to be used by the language model, BLIP adapts visual tokens to be ingested just like text tokens by the language model. This idea has been further developed in more recent papers [12, 9, 32, 54]. In this work, we investigate if instead of representing the image by latent visual features, we can directly provide textual representations of an image by focusing on its elements: objects and regions.
We explore this alternative strategy with the idea of flexible captioning – generating controllably rich and localized captions as shown in Figure 1. Our model, FlexCap, enables spatially controllable inquiry of any bounding box in the image, with the desired text detail controlled by the generated word count. FlexCap effectively combines three tasks that have been studied in isolation until now: image captioning, object detection, and dense captioning. While image captioning models can capture coarse semantic information they lack spatial understanding of the visual content. On the other hand, object detection brings spatial information in the form of bounding boxes. But object detection systems lack semantic details (attributes and relationships between objects) and are usually limited to a few classes. Recently open-vocabulary object detection methods [35, 15] increase the semantic diversity but it is still limited by text queries or prompts provided manually.
Dense captioning [25] is the task of localizing salient regions of the image and describing them with natural language sentences. However the capacity and richness of these models are limited with existing image captioning datasets (e.g. COCO[31] and Visual Genome [27]). Moreover, these methods focus on full sentence generation rather than a mix of short and long captions at several levels of richness as people typically do. FlexCap combines all three tasks into one system by formulating each of them as different captioning problems – image captioning implies using FlexCap to caption the whole image as one big bounding box, dense captioning can be performed by conditioning on individual boxes, and object detection can be performed by prompting the model to produce short class names as captions.
@@
To be able to train such a model, we require a dataset of images where many boxes are labeled with short and long descriptions. We propose a method to generate triplets of (i) image, (ii) a proposed region within the image, and (iii) its corresponding caption from image-text pair datasets, through the use of open-vocabulary object detectors. We demonstrate this at two scales: 200 million triplets using YFCC100M [46] captioned images; and 32 billion triplets using the WebLI [9] captioned dataset. Training FlexCap on this dataset enables the model to generate spatially and semantically rich representations (bounding boxes and their descriptions) that focuses on objects, their attributes, and their contextually changing descriptions. We show that this human interpretable representation, when combined with the power of LLMs, enables visual question answering and dialog. We also demonstrate that this combination can result in performance that is competitive with state-of-the-art VLM models on zero-shot image and video question answering benchmarks.
@@
Our key technical contributions: (i) controllable localized visual descriptions, using word count as a proxy for complexity to modulate the output of a generative language model, and using bounding box conditioning to indicate local regions in the image (see Sec 2) (ii) a large-scale dataset generation method for producing image-box-text triplets required to train our model. (see Sec 3) (iii) structured and human-interpretable visual representations that can be input to LLMs and achieve comparable or state-of-the-art performance in VQA tasks. (see Sec 4 and Sec 5.1). (iv) demonstrating that FlexCap enables better localized captioning and open-ended object detection than baselines under comparable scenarios. (see Sec 5.2 and Sec 5.3). (v) highlighting the emergent property of FlexCap to extract diverse visual information using prefix conditioning. (see Sec 5.4)
@@ 2 FlexCap
Architecture. Our objective is to train a model that takes as input an image and a region of interest and outputs a description of the region spanned by the box. We present FlexCap’s architecture in Figure 2. The model takes an image, the coordinates of a bounding box and the conditioning tokens as input, and outputs a textual description of visual contents within the specified bounding box. Our model mainly consists of an image encoder (i.e. SOViT-400M/14 [1]) and a transformer-based text-decoder. We pass the image through the vision model to produce outputs of dimensions $n\times d$ (where $n$ is the number of patches and $d$ is the embedding size). We pass the bounding box coordinates (of dimensions $1\times 4$) through a linear layer to produce the coordinate features (of dimension $1\times d$). Both vision features and normalized bounding box features are concatenated to form input of dimension $(n+1)\times d$ to a text decoder. The text decoder consists of a stack of $L$ Transformer layers. We use a decoder-only architecture in which all the vision and bounding box tokens remain unmasked but the text tokens are masked in a causal manner to enable next-word prediction training. We add all the vision tokens and bounding box coordinate tokens so that the text decoder can access all of the visual context in the image and the exact bounding box location. In this work, we train a text decoder composed of 12 self-attention transformer layers with a dimensionality of 768 and 12 attention heads.
@@
Length conditioning. For the same region in an image, there may be multiple valid captions. In the input image shown in Fig. 3, all the following descriptions are correct: dog, Border collie, dog playing with a frisbee, brown dog playing on a grass field. Clearly, the task of describing a bounding box in the image does not have only one right answer. We equip the model with the capability of producing outputs of a desired length by utilizing the idea of length conditioning. We condition the input by simply using an additional token indicating the desired length of the output caption. Training with length-conditioning is useful for three key reasons. First, the number of words used to describe is often proportional to the information content. We train the model to predict the next words in the sequence while accounting for the desired length, thereby the model learns to modulate the amount of information in the generated text. Second, length conditioning allows users to control the output of the model further enabling the use of a single model for many diverse tasks. Third, the length prefix provides a better conditioned initial state for the captioner. Fig. 2 shows how the same box might have more than one ground truth captions <s> a cat <e> or <s> grey and white cat lying on shoes <e>. If we use the first caption as ground truth and the words <s> a cat as the seen text, the next-word prediction loss encourages the model to increase its score for the <e> token and decrease the score for the word playing due to the softmax loss. This is in direct conflict with the second caption. We can alleviate such training issues by using length prefix for text generation.
@@
Loss. We train FlexCap to predict the next token of the text. The text tokens are prefixed with the desired length of the caption and appended with an end of sentence token <e> to indicate the end of the caption. The target text tokens are obtained by shifting the padded text by 1. This is common training methodology for training generative language models like GPT [7] or SimVLM [47]. The loss is a classification loss over all the words present in the vocabulary. The loss is ignored over the padded tokens that are used to keep the size of the outputs same for all the captions in the batch.
Formally, we describe our loss formulation below. We represent a data sample as a triplet $T=(X,B,W)$ consisting of image $X$, bounding box $B$ and captions $W$, where $W=\{\text{LENGTH-K},w_{1},w_{2},...w_{k}\}$. To enable batch training, we pad the tokenized captions to a fixed size $M$. For a given data triplet, our objective is to maximize the following log-likelihood. $l(X,B,W)=\sum^{M}_{i=1}\log p(w_{i}|w_{<i},X,B)$. Assume that we have a dataset $\mathcal{D}=\{T_{1},T_{2},...,T_{N}\}$. The overall loss function is:
@@
Implementation. We implement this model using the JAX framework [6]. We train the entire model for about $400K$ steps using the AdamW optimizer with a cosine learning rate schedule. The maximum learning rate is $1.6\times 10^{-4}$ with 10K warm-up steps. We use a weight decay of $0.05$. We train with a batch size of $4096$ and image resolution of $224\times 224$. We use a maximum text sequence length of $32$ tokens. For each image in the batch, we sample a maximum of 8 bounding boxes for each training step.
Inference. At inference time, we provide an image, the target bounding box, and the desired length as input. We then decode in an auto-regressive manner till the end of caption token <e> is encountered or the maximum number of decoding steps is reached. While we can use standard sampling techniques used in text-generation like beam search, temperature sampling, or nucleus sampling [17] to generate multiple captions, we use greedy decoding in all experiments for simplicity.
@@ 3 Localized Captions Dataset
In order to train the FlexCap model, we build a large scale dataset of image region descriptions of varying lengths. In the following section we describe how we produce such a dataset from existing image-text paired datasets. We leverage the web-based image-caption pairs datasets (like WebLI [9] and YFCC100M [46]) to create localized captions dataset. The dataset generation pipeline is shown in Fig. 3. First we create text queries using n-grams from the caption of the image: e.g. “dog",“brown dog", “brown dog playing with a disc". We specifically create n-grams where $n=\{1,2,\cdots,8\}$ and then filter out incomplete captions like “with a red", “dog playing with". More details about the filtering step are mentioned in the appendix. Then we use the filtered n-grams as text queries for pre-trained region proposal models (i.e.  OWL-ViT [36]) to extract boxes and select text-box pairs based on the similarity score ($>$ 0.1). Multiple n-grams may match for a box, and this results in several ways of describing a box in the image as shown in Col. 4 in Fig. 3.
@@
YFCC100M. We also create a localized captions dataset using YFCC100M images. Specifically we use the same 14M images as the CLIP paper. The dataset creation method results in $\sim 11$M images with at least one valid box. On average each image has $20$ boxes, making the size of the localized captions YFCC100M dataset is $\sim 0.2$B image-box-caption triplets. The number of YFCC100M triplets is  160 times smaller than the localized captions dataset created from WebLI.
As both OWL-ViT and the CLIP subset of YFCC100M are publicly available, the resulting localized captions dataset can be generated with open-source models and public datasets. Since the WebLI dataset is not publicly available yet, YFCC100M triplet-dataset can serve as a reproducible benchmark.
@@
4 FlexCap with LLMs
Recent work on improving the recognition and reasoning capability of vision models, so that they can perform tasks such as question answering and visual dialog, has focused on connecting features from vision models like ViT to Large Language Models (LLMs). Existing works like Flamingo [3], LlaVA [32], PALI [9], PALM-E [12], Ferret [54], LaVIT [24], have focused on using patch features from images as words or tokens. Here, we propose an alternate approach.
Instead of adapting LLMs to understand vision features, we explore connecting images with LLMs using rich text in the form of localized descriptions. As FlexCap solves the rich visual perception problem by generating dense localized information, it enables LLMs to perform high-level spatial and textual reasoning by linking spatially enriched visual concepts with the world knowledge. When combined with their well-trained common-sense reasoning, LLMs can reason about attributes by looking at singular object detections, scene-level understanding by utilizing larger window descriptions, counts by checking number of bounding boxes with similar content, and relative spatial relationships by comparing location of the boxes. While using captions as input for LLMs is not novel in itself [40, 48], ensuring the captions are diverse, descriptive, and exhaustively describe the full image is challenging. Now we will discuss how to connect FlexCap with LLMs.
FlexCap-LLM. In Figure 4, we show how we use FlexCap with an LLM to solve visual questions. First, we convert an image to a sequence of localized descriptions that describe the image in terms of the objects and regions present in the image. To do so, we need region proposals. We use OWL-ViTv2 [34] to localize important objects and regions in an image. We keep the top 128 bounding boxes by their objectness scores. We then use FlexCap to describe each box in the image in the context of the entire image. In order to produce holistic descriptions, we use multiple prefixes for each region. These prefixes are a combination of length conditioning token and some initial text. We add the boxes and their descriptions to a text preamble (see Fig. 4) that defines the setup where we are using an LLM to answer questions about an image. In all the experiments, we use PALM2-S model [4] as the LLM of choice. We refer to this end-to-end system that takes an image and a question to output the answer as FlexCap-LLM.
@@ 5 Experiments
We evaluate our model across different tasks that require different levels of detail. In Section 5.1, we test FlexCap-LLM on visual question answering tasks. In Section 5.2, we evaluate how FlexCap can be used for localized captioning. In Section 5.3, we compare the effectiveness of using FlexCap and other VLMs for the task of open-ended object detection. In Section 5.4, we present qualitative results and applications of FlexCap. In Section 4, we provide ablations and analysis of the generated captions.

### v2

@@ Abstract
We introduce FlexCap, a vision-language model that generates region-specific descriptions of varying lengths. FlexCap is trained to produce length-conditioned captions for input boxes, enabling control over information density, with descriptions ranging from concise object labels to detailed captions. To achieve this, we create large-scale training datasets of image region descriptions with varying lengths from captioned web images. We demonstrate FlexCap’s effectiveness in several applications: first, it achieves strong performance in dense captioning tasks on the Visual Genome dataset. Second, we show how FlexCap’s localized descriptions can serve as input to a large language model to create a visual question answering (VQA) system, achieving state-of-the-art zero-shot performance on multiple VQA benchmarks. Our experiments illustrate FlexCap’s utility for tasks including image labeling, object attribute recognition, and visual dialog. Project webpage: https://flex-cap.github.io.
@@ 1 Introduction
How does one describe the world around us, not just in broad strokes but with the ability to zoom in and out, capturing both the grand scene and the minute details? Imagine pointing at a bustling market scene and asking, "What’s happening here?" and receiving a vivid description, not just of the market as a whole, but also a detailed account of the interactions between vendors and customers, the vibrant colors of the goods on display, or even a specific item held by a passerby. This ability to controllably focus and describe visual content is what we call flexible captioning.
Traditional image captioning models, while adept at capturing the gist of an image, often struggle to pinpoint specific objects or attributes. On the other hand, object detection systems excel at localizing elements but may lack the vocabulary to describe them comprehensively. Dense captioning [26] attempts to bridge this gap by generating captions for multiple regions, but its expressiveness is limited by existing datasets. In this work, we introduce a model called FlexCap that bridges the gap between holistic image understanding and localized inquiry. This enables the generation of captions that are both spatially precise and semantically rich (see Fig. 1 (left)) by specifying a region of interest and the desired level of detail in terms of number of words in the predicted caption. This allows us to integrate the strengths of image captioning, object detection, and dense captioning into one model.
@@
To be able to train such a model, we require a dataset of images where many boxes are labeled with short and long descriptions. We propose a method to generate triplets of (i) image, (ii) a proposed region within the image, and (iii) a caption of a particular length, by using an open-vocabulary object detector to label regions from captions of an image-text pair dataset. We demonstrate this at two scales: 200 million triplets using YFCC100M [47] captioned images; and 32 billion triplets using the WebLI [9] captioned dataset. Training FlexCap on these datasets enables the model to generate spatially and semantically rich representations (bounding boxes and their descriptions) that focuses on objects, their attributes, and their contextually changing descriptions.
We evaluate the effectiveness of the descriptions generated with FlexCap by representing images in detail by describing its constituent objects and regions. A popular and effective strategy is to directly provide visual features as input tokens to LLMs  [2, 30, 29, 14, 9, 33, 59]. Instead we directly provide textual descriptions of images as input to the LLM to evaluate FlexCap similar to [57, 21, 42]. We show that this human interpretable representation, when combined with the power of LLMs, enables visual question answering and dialog. An example is shown in Figure 1(right). We also demonstrate that this combination can result in performance that is competitive with state-of-the-art VLM models on zero-shot image and video question answering benchmarks.
@@
Our key technical contributions are: (i) controllable localized visual descriptions, using word count as a proxy for complexity to modulate the output of a generative language model (ii) a large-scale dataset generated for image-text-box triplets to enable training of our model; (iii) the human-interpretable representations produced by FlexCap with the help of LLMs is comparable or exceeds performance of state-of-the-art VQA methods; and (iv) demonstrating that our dense captioning methods outperform baselines under comparable scenarios.
@@ 2 FlexCap
Length conditioning. For the same region in an image, there may be multiple valid captions. In the input image shown in Figure 2, all the following descriptions are correct: cat, grey and white cat, grey and white cat lying on shoes. Clearly, the task of describing a bounding box in the image does not have only one right answer. We equip the model with the capability of producing outputs of a desired length by utilizing the idea of length conditioning. We condition the input by simply using an additional token indicating the desired length of the output caption. Training with length-conditioning is useful for three key reasons. First, the number of words used to describe is often proportional to the information content. We train the model to predict the next word in the sequence while accounting for the desired length, thereby the model learns to modulate the amount of information in the generated text. Second, length conditioning allows users to control the output of the model, further enabling the use of a single model for many diverse tasks. Third, the length prefix provides a better conditioned initial state for the captioner. Figure 3 shows how the same box might have more than one ground truth caption <s> a dog <e> or <s> dog playing with a frisbee <e>. If we use the first caption as ground truth and the words <s> a dog as the seen text, the next-word prediction loss encourages the model to increase its score for the <e> token and decrease the score for the word playing due to the softmax loss. Whether this occurrence will be a problem depends on the dataset statistics. To quantify the prevalence of this problem, we compute a statistic: for each image box, we consider all pairs of captions and measure the fraction of pairs sharing prefix words. For instance, consider a box that has three captions: <s> dog <e>, <s> dog playing <e>, and <s> dog playing with a frisbee <e> which share a prefix <s> dog. While another caption for the same box, <s> brown dog <e>, does not share a prefix with captions beginning with <s> dog <e>. After averaging this metric across all images in our Localized Captions Dataset from WebLI (introduced later in Section 3), we found that 30.8% of all caption pairs share a prefix. Using a length conditioning token instead of <s>, the probability of prefix matching decreases from 30.8% to 11.1%. The length conditioning helps the model in distinguishing between captions with the same prefix while also providing the model with a novel capability during inference.
Architecture. Our objective is to train a model that takes as input an image and a region of interest and outputs a description of a desired length of the region spanned by the box. We present FlexCap’s architecture in Figure 2. The model takes an image, the coordinates of a bounding box and the conditioning tokens as input, and outputs a textual description of visual contents within the specified bounding box. Our model mainly consists of an image encoder (i.e. SOViT-400M/14 [1]) and a transformer-based text-decoder. We pass the image through the vision model to produce outputs of dimensions $n\times d$ (where $n$ is the number of patches and $d$ is the embedding size). We pass the bounding box coordinates (of dimensions $1\times 4$) through a linear layer to produce the coordinate features (of dimension $1\times d$). Both vision features and normalized bounding box features are concatenated to form input of dimension $(n+1)\times d$ to a text decoder. The text decoder consists of a stack of $L$ Transformer layers. We use a decoder-only architecture in which all the vision and bounding box tokens remain unmasked but the text tokens are masked in a causal manner to enable next-word prediction training. We add all the vision tokens and bounding box coordinate tokens so that the text decoder can access all of the visual context in the image and the exact bounding box location. In this work, we train a text decoder composed of 12 self-attention transformer layers with a dimensionality of 768 and 12 attention heads.
@@
Loss. We train FlexCap to predict the next token of the text. The text tokens are prefixed with the desired length of the caption and appended with an end of sentence token <e> to indicate the end of the caption. The target text tokens are obtained by shifting the padded text by 1. This is common training methodology for training generative language models like GPT [8] or SimVLM [51]. The loss is a classification loss over all the words present in the vocabulary. The loss is ignored over the padded tokens that are used to keep the size of the outputs same for all the captions in the batch. Formally, we represent a data sample as a triplet $T=(X,B,W)$ consisting of image $X$, bounding box $B$ and captions $W$, where $W=\{\text{LENGTH-K},w_{1},w_{2},...w_{k}\}$. To enable batch training, we pad the tokenized captions to a fixed size $M$. For a given data triplet, our objective is to maximize the following log-likelihood. $l(X,B,W)=\sum^{M}_{i=1}\log p(w_{i}|w_{<i},X,B)$. Assume that we have a dataset $\mathcal{D}=\{T_{1},T_{2},...,T_{N}\}$. The overall loss function is:
@@
Implementation. We implement this model using the JAX framework [7]. We train the entire model for about $400K$ steps using the AdamW optimizer with a cosine learning rate schedule. The maximum learning rate is $1.6\times 10^{-4}$ with 10K warm-up steps. We use a weight decay of $0.05$. We train with a batch size of $4096$ and image resolution of $224\times 224$. We use a maximum text sequence length of $32$ tokens. For each image in the batch, we sample a maximum of 8 bounding boxes.
Inference. At inference time, we provide an image, the target bounding box, and the desired length as input. We then decode in an auto-regressive manner till the end of caption token <e> is encountered or the maximum number of decoding steps is reached. We use greedy decoding in all experiments unless otherwise stated. However, we can use standard sampling techniques used in text-generation like beam search, temperature sampling, or nucleus sampling [19] to generate multiple captions.
@@ 3 Localized Captions Dataset
In order to train the FlexCap model, we build a large scale dataset of image region descriptions of varying lengths. In the following section we describe how we produce such a dataset from existing image-text paired datasets. We leverage the web-based image-caption pairs datasets (like WebLI [9] and YFCC100M [47]) to create a localized captions dataset. The dataset generation pipeline is shown in Figure 3. First, we create text queries using n-grams from the caption of the image: e.g. “dog",“brown dog", “brown dog playing with a disc". We specifically create n-grams where $n=\{1,2,\cdots,8\}$ and then filter out incomplete captions like “with a red", “dog playing with". More details about the filtering step are mentioned in the appendix. Then we use the filtered n-grams as text queries for pre-trained region proposal models (i.e.  OWL-ViT [36]) to extract boxes and select text-box pairs based on the similarity score ($>$ 0.1). Multiple n-grams may match for a box, and this results in several ways of describing a box in the image as shown in Col. 4 in Figure 3.
@@
YFCC100M. We also create a localized captions dataset using YFCC100M images. Specifically we use the same 14M images as the CLIP paper. The dataset creation method results in $\sim 11$M images with at least one valid box. On average each image has $20$ boxes, making the size of this dataset $\sim 0.2$B image-box-caption triplets. The number of YFCC100M triplets is  160 times smaller than the localized captions dataset created from WebLI.
As both OWL-ViT and the CLIP subset of YFCC100M are publicly available, the resulting localized captions dataset can be generated with open-source models and public datasets. Since the WebLI dataset is not publicly available yet, YFCC100M triplet-dataset can serve as a reproducible benchmark. Concurrently large-scale grounded image-text dataset generation pipelines have also been proposed in [38] and  [50].
@@ 4 Experiments
Figure 4: Evaluating open-vocabulary outputs from FlexCap using the CLIP [39] text encoder.
4.1 Correctness and Compliance of Generated Captions
Correctness. In this experiment, we solely evaluate the recognition capabilities of our model in a zero-shot manner. We use the region classification task [53, 61] on the MS-COCO dataset to assess how well our model recognizes objects at different scales and under occlusion. In this task, the image and the ground truth bounding box are provided as input to the model such that it produces a short description of what is contained in the bounding box. Our region classification pipeline (Figure 4) passes the input image (of size $448\times 448$) through FlexCap which generates 20 captions each for 4 caption lengths (1, 2, 3, 4) via nucleus sampling. These captions are then mapped to object class names using CLIP’s [39] text encoder. By comparing the mean of the text embeddings of predicted captions with those of ground truth class names, we obtain the classification scores. We report the results in Table 1(a). We find FlexCap outperforms contrastively trained approaches used for region classification. Furthermore, we observe a significant boost ( 13% mAP) in performance by producing multiple captions for each box. One reason for the boost in performance may be directly comparing text embeddings to produce classification scores as compared to baselines which use dot product of image and text embeddings.
Compliance. In Figure 5, we show qualitative examples of the FlexCap model producing different length captions for the same box. Note how the model progressively adds more information about the object by incorporating context in the longer sentences (in the jungle), attributes (pink flamingo kite), and alternative nouns (chevy, feline). We also measure how well our model complies to the desired caption length. To do so, we take 1000 images from the MS-COCO dataset and use a random object in the image to produce descriptions with different lengths. We report the average length of the predicted caption, and the fraction of times the predicted caption has a length equal to the target length in Table 1(b). We find FlexCap’s outputs are mostly compliant with the target length.
Method mAP ($\uparrow$)
CLIP [53] 29.2
CLIM [53] 62.2
RegionCLIP [61] 62.8
FlexCap (Top-1 caption) 72.0
FlexCap (Top-20 captions) 85.0
(a) MS-COCO Region Classification with ground truth bounding box.
Target Mean Len Accu- Target Mean Len Accu-
Length of Pred. racy Length of Pred. racy
1 1.02 0.99 5 5.06 0.94
2 2.05 0.96 6 6.01 0.97
3 3.06 0.95 7 7.02 0.95
4 4.04 0.96 8 8.02 0.95
(b) Compliance metrics. FlexCap produces length-compliant captions for different lengths.
Table 1: FlexCap’s outputs are accurate and length compliant.